If you bought 850 Douglas firs and ponderosa pines and paid an average of $300 for douglas firs and $225 for ponderosa pines and the total you spent was $217,500 how many of each tree did you buy?
Answer by HyperBrain(694)

Let x= no. of firs
y= no. of ponderosa
(1)x+y=850
(2)300x+225y=217500
Substituting (1) to (2),
300*(850-y)+225y=217500
255000-300y+225y=217500
255000-75y=217500
75y=3750
y=500
x+500=850
x=350
Therefore, he bought 350 firs and 500 ponderosas!